The balance sheet and income statement for Covid19 Bhd is presented below.

BALANCE SHEET (000)

Cash $   500

Accounts receivable 1,500

Inventories                                                  500

Current assets 2,500

Net fixed assets                                        5,000

Total Assets $ 7,500

Accounts payable                                     1,200

Bank note                                                    300

Total current liabilities 1,500

Long term debt 4,000

Common stock                                            300

Retained earnings                                     1,700

Total liabilities and owner's equity $ 7,500

INCOME STATEMENT (000)

Net sales                                                 $ 8,500

Cost of goods sold                                   (3,400)

Gross profit 5,100

Operating expenses                                 (2,900)

Net operating income 2,200

Interest expense                                         (580)

Earnings before taxes 1,620

Income tax (34%) (551)

Net income $ 1,069

Answer the following questions:

  1. Compute the following ratios: Current ratio, Acid test ratio, Debt ratio, Total asset turnover, Operating profit margin, Net profit margin, Times interest earned, Inventory turnover.
  2. All other things equal, compute the dollar amount of sales needed to achieve an 18% return on total assets for the coming year.
Answer #1

Ans:

(a)

1. Current Ratio = Current Assets/Current Liabilities

= 2500/1500

=1.67

2. Acid Test Ratio/Quick Ratio = [Current Assets – Inventory – Prepaid Expenses] / Current Liabilities

= (2500-500-0)/1500

=2000/1500

=1.33

3. Debt Ratio = Total Liabilities(long term and short term debt)/Total Assets

=4000/7500

=53.33%

4. Total assets Turnover Ratio = Net Revenue or Net sales/ Average total assets

=8500/7500

=1.13

5. Operating Profit Margin Ratio = Net operating Income or EBIT/Total sales or revenue

= 2200/8500 * 100

= 25.88%

6. Net Profit Margin Ratio = Net profit/Revenue

= 1069/8500

= 12.58%

7. Times Interest earned Ratio = Earning before Interest and tax (EBIT)/ Total interest payable on bonds and other debt

= 2200/580

= 3.80

8. Inventory Turnover Ratio = 1/ Inventory days ratio ÷ 265 (see working Note)

= 1/ 54 ÷ 365

=1/0.15

= 6.75 approx

Working note-

Inventory days ratio = (Inventory / Cost of goods sold ) * 365

= (500/3400) * 365

= 54 days approximately

(b) Return on total assets = Earning after tax (EAT) / Total assets

18% = EAT/7500

EAT = 1350

Sales needed to achieve an 18% return on total assets for the coming year:

Earnings after tax (EAT) $1350
Add: Taxes ($1350 / 66) * 34 $696
Earnings before tax (EBT) $2046
Add: Interest $580
EBIT $2626
Add: Operating expenses $2900
Net operating income $5526
Add: Cost of goods sold $3400
Net Sales $8926